OpenStudy (anonymous):

Help!!!:The ratio of apples to oranges at a fruit stand Is 4 to 5.There are 60 oranges at the stand. How may apples are at the fruit stand? A. 48 B. 50 C.59 D. 75

OpenStudy (anonymous):

@Jadeishere @viop9

OpenStudy (jadeishere):

4/5 = 60/x 4x = 300 divide 300 by 4 and you have your answer